OpenStudy (anonymous):

Solve graphically: 3y = 5 - x and 2x = y + 3. Write down the co-ordinates of the point of intersection of the two lines.

OpenStudy (anonymous):

rewrite slope standard from 3y+x=5 time (-2) u get -6y-2x=-10 - y +2x =3 -y +2x = 3 Add ----------------- - 7y = -7 y = -7/-7 y = 1 Sutitution y =1 solve for x 3(1) = 5 - x 3 = 5 -x 3-5 =-x -2/-1= x 2 = x

OpenStudy (anonymous):

Thank you.